# Genomic Language Models: Opportunities and Challenges

### v1

@@ INTRODUCTION
The sequence embeddings extracted from language models are commonly used as representations that capture rich contextual information and sequence features. Unsupervised clustering of GPN’s final-layer embeddings shows distinct clusters of input sequences that correspond to different genomic classes such as CDS, intronic, UTR, etc. [13] (Figure 2d). Similarly, unsupervised clustering of SpliceBERT embeddings of canonical splice sites and non-splice GT/AG sites reveals distinct clusters that correspond to the two groups, implying that the model has learned to capture key contextual patterns that determine functional elements in the genome [52].
The attention mechanism in the Transformer model is designed to capture the pattern of interaction between input tokens. Thus, interpreting the attention weights or the attention map for a given input sequence can reveal genomic features learned by the model. In SpliceBERT, attention weights between splice donors and acceptors are significantly higher than those between random pairs of sites; also, the strength of interaction tends to be higher within true donor-acceptor pairs compared to other combinations of donor and acceptor sites. These findings suggest that the model has learned the relationship between functionally interacting sites [52].
@@
Here, we discuss how gLMs can be evaluated in regards to the three application areas described earlier: predicting alleles’ fitness, generating novel viable sequences, and transfer learning.
One way to benchmark a fitness predictor is against experimentally obtained measurements from multiplexed assays of variant effect [93, 94, 95]. These assays couple functional differences between genetic variants to readouts (such as the expression of a reporter gene). In turn, these readouts may be used to rank variants by their functionality. Since variants that affect function also tend to affect fitness, we should expect that experimentally obtained and predicted ranks of variants should be correlated. One source for these data is ProteinGym, a widely-used collection of experimental data that may be used to benchmark missense variant effect predictors [96]. Another label that may reflect fitness is whether variants have evidence of pathogenicity – that is, can elevate the risk for diseases. Pathogenic variants may affect fecundity, and, therefore, be deleterious. As a result, we can benchmark fitness predictors by evaluating them as pathogenicity classifiers. In human genetics, primary sources of clinical labels for variants include the ClinVar [97], HGMD [98], and OMIM [99] databases. A third type of data that provides information on alleles’ fitness values are variant frequencies. Since common variants are unlikely to be highly deleterious [100], their predicted fitness values generally should be relatively higher than those of rare variants. Therefore, we may benchmark predictors based on how well they identify common variants. A primary source of data on human allele frequencies in various ancestry groups is the gnomAD database [101].
@@
The murky relationship between these data and fitness compounds the problem of extrapolating predictors’ operational performance from their performance on benchmarks. Predictors may appear to excel, but may do so by exploiting the ways in which benchmarks fail to capture fitness. For example, a critical issue with using clinical labels is that variants are classified based on whether there is ample evidence that they are benign or pathogenic [106]. Since predictors may also utilize this evidence, their benchmarked performance on labeled variants may not reflect their true performance on unlabeled variants. (See Box 3 for a brief discussion of generalization performance.) There are also critical issues with using allele frequency data: for one, in addition to the direct action of natural selection, allele frequencies are influenced by factors such as mutation rates, drift, background selection, and hitchhiking [107]. As a result, predictors may perform well on benchmarks by predicting the effects of these processes instead of fitness. These issues highlight a need to carefully interpret the causes of predictors’ performance, and they have led to calls for greater transparency on which data and methods are used to train predictors [108].
@@
Regarding sequence design evaluation, one should adopt a holistic approach and examine a broad range of properties of the generated sequences. For instance, Polygraph [109], a recent benchmark for regulatory sequence design, proposes a series of analyses that investigate the sequence composition, motif pattern, and predicted functional activity. In whole-genome or chromosome design tasks, one should further evaluate the existence and positioning of essential genes and functional regulatory elements, as well as the interactions between them. Ultimately, the designed sequences should be evaluated via in vivo or in vitro experiments to determine if they perform the desired function.
The problem of evaluating the utility of gLMs for transfer learning is uniquely challenging, as pretrained models are intended to be applied to a broad and perhaps indeterminate set of tasks. A solution is to evaluate models on a representative subset of tasks and extrapolate their broader utility. One approach is to process data from representative functional genomics experiments (such as those from the ENCODE [110] or Roadmap Epigenomics [111] projects) into genomic and variant annotations that models can be trained to predict from their sequence context. To facilitate comparison between models, these annotations have been consolidated into various standardized sets of training and test data [16, 112, 60, 39]. However, a problem with these benchmarks is that they are often either opaque and possibly say little about the adaptability of models, or they are relatively trivial, evaluating models against annotations that are themselves derived from simple computational methods (which may themselves be heuristic and are surely wrong some of the time). This makes it difficult to tell whether models that perform better on these benchmarks are genuinely more adaptable, and whether outstanding performance on these benchmarks is sufficient for a model to be broadly adaptable.
@@
[style=box]
List of Boxes 4 Outstanding Questions for Future Research
@@
    Will the scaling hypothesis hold for gLMs, and for how long? Are there really that much data available, considering that most may be non-functional? If we do need scale, how can we organize funding and compute resources to ensure that models can be trained by academic labs and not solely depend on a few industry players with unclear incentives?
CONCLUDING REMARKS & FUTURE PERSPECTIVES
In an age of a vast and growing number of genomic sequences, gLMs are emerging as powerful tools to extract complex patterns useful for numerous applications, including fitness estimation, sequence design and transfer learning. However, they do not yet represent a magical, sudden breakthrough as the term “AI” may suggest. Instead, we view them as another useful modeling tool, much like Hidden Markov Models were when they were first introduced.
As we introduce new techniques into the field of biology, we can provide more value by using clear vocabulary and drawing connections to existing concepts rather than introducing unnecessary jargon that exaggerates novelty (which can confuse readers, editors and funding sources). While the distinction is occasionally useful, “self-supervision” is in many aspects just another unsupervised learning technique. Similarly, in most biological tasks “zero-shot” prediction could be rephrased as unsupervised prediction. Finally, designating a pretrained model as “foundation” or “frontier” gives the impression that the model has undeniably redefined a field. This might be applicable to GPT [113], but hardly to current generation models in genomics, which are often insufficiently benchmarked to begin with.
While earlier gLMs tend to be more or less direct adaptations from NLP models, we expect that further contextualization with deep genomics expertise will reap the highest rewards. We note that evaluating the capabilities of gLMs is challenging because metrics may be misleading, especially when over-optimized. A boon for NLP is that humans are experts in natural language and, therefore, can calibrate benchmarks to match their expertise. In genomics, however, we must rely on data and expert knowledge to falsify models. This aspect of the problem makes it especially challenging and may highlight a need for engagement with subject-matter experts and deliberate experimentation for the sake of developing benchmarks. We conclude this review with a set of outstanding questions in Box 4 that we believe warrant further investigation.

### v2

@@ INTRODUCTION
The sequence embeddings extracted from language models are commonly used as representations that capture rich contextual information and sequence features. Unsupervised clustering of the encoded sequence embeddings from gLMs has shown distinct clusters of input sequences that correspond to different genomic classes such as CDS, intronic, UTR, etc. [54, 13, 16, 35] (Figure 2d). Additionally, unsupervised clustering of SpliceBERT embeddings of canonical splice sites and non-splice GT/AG sites reveals distinct clusters that correspond to the two groups [64]. These results suggest that the models have learned to capture key contextual patterns that characterize functional elements in the genome.
The attention mechanism in the Transformer model is designed to capture the pattern of interaction between input tokens. Thus, interpreting the attention weights or the attention map for a given input sequence can reveal genomic features learned by the model. In SpliceBERT [64], attention weights between splice donors and acceptors are significantly higher than those between random pairs of sites; also, the strength of interaction tends to be higher within true donor-acceptor pairs compared to other combinations of donor and acceptor sites. These findings suggest that the model has learned the relationship between functionally interacting sites.
@@
More recently, the dependency between genomic positions learned by a gLM was studied by introducing point mutations at a position and quantifying the changes in nucleotide probabilities at other positions [19]. Nucleotide dependency analysis revealed learned interactions within and across functional elements such as TFBS, splice sites and RNA, including known secondary and tertiary structure contacts. Notably, nucleotide dependency analysis was able to detect bound TFBS more robustly than the previous approach based on predicted marginal probability distributions.
@@
In this section, we discuss how models’ performance can be benchmarked in regards to the three application areas described earlier: predicting functional constraints on alleles, generating novel viable sequences, and transfer learning.
There are various types of data that reflect functional constraint on alleles and may be used to benchmark a variant effect predictor. One type of data are assays that couple functional differences between genetic variants to readouts (such as the expression of a reporter gene or cell growth) [110, 111, 112]. These readouts may be used to rank variants by their functionality, and since variants that affect function also tend to be under selection, we should expect that these ranks should correlate with ranks obtained from models’ predictions. One source for these data is ProteinGym, a widely-used collection of experimental data that may be used to benchmark missense variant effect predictors [113]. Another type of data are clinical labels indicating whether variants have evidence of pathogenicity—that is, can elevate the risk for diseases. Pathogenic variants may affect fecundity, and, therefore, be deleterious. As a result, we can benchmark variant effect predictors by evaluating them as pathogenicity classifiers. In human genetics, primary sources of clinical labels for variants include the ClinVar [114], HGMD [115], and OMIM [116] databases. A third type of data are variant frequencies. Since common variants are unlikely to be highly deleterious [117], their predicted level of constraint should be relatively higher than those of rare variants. Therefore, we may benchmark predictors based on how well they identify common variants. A primary source of data on human allele frequencies in various ancestry groups is the gnomAD database [118]. Altogether, these data may be used as separate lines of evidence for models’ generalization performance.
@@
An issue with the evaluation of variant effect predictors is that the relationship between validation data and functional constraint may be murky. As a consequence, models may excel at benchmarks by exploiting the ways in which data fail to capture functional constraint. For example, a critical issue with using clinical labels is that variants are classified based on whether there is ample evidence that they are benign or pathogenic [123]. Since predictors may also utilize this evidence, their benchmarked performance on labeled variants may not reflect their true performance on unlabeled variants. (See Box 3 for a brief discussion of generalization performance.) There are also critical issues with using allele frequency data: for one, in addition to the direct action of natural selection, allele frequencies are influenced by factors such as mutation rates, drift, background selection, and hitchhiking [124]. As a result, predictors may perform well on benchmarks by predicting the effects of these processes instead of functional constraint. These issues highlight a need to carefully interpret the causes of predictors’ performance, and they have led to calls for greater transparency on which data and methods are used to train predictors [125].
@@
There are a separate set of challenges with the evaluation of generative sequence models. A basic way to evaluate language models’ generative capabilities is to compare their perplexities on sets of valid sequences. However, to evaluate models’ capability to design novel sequences, it is necessary to gauge whether they can identify sequences that are both viable and novel. For this reason, models’ perplexities on test sets may not reliably indicate their utility for design. Instead, a holistic approach that examines a broad range of properties of generated sequences may be warranted. For instance, Polygraph [126], a recent benchmark for regulatory sequence design, proposes a series of analyses that investigate sequence composition, motif patterns, and predicted functional activity. For whole-genome or chromosome design tasks, it may also be necessary to evaluate the existence and positioning of essential genes and functional regulatory elements, as well as the interactions between them. Ultimately, the designed sequences should be experimentally evaluated to determine if they perform their desired functions.
Lastly, there is a unique challenge with the evaluation of gLMs for transfer learning: any set of benchmarks—perhaps, in conjunction—must reliably indicate a model’s performance on relevant tasks. A type of data that may be fashioned into a set of tasks broadly informative of models’ adaptability to genome interpretation are functional genomics data (such as those from the ENCODE [127] or Roadmap Epigenomics [128] projects), which may be used to annotate genomic regions and variants. We should expect that a model’s performance on predicting these annotations from genome sequences after adaptation is indicative of their capability to identify functionally similar genomic elements. To facilitate comparison between models, these annotations have been consolidated into various standardized sets of training and test data [16, 129, 71, 50].
As transfer learning benchmarks help highlight limitations of current models and establish criteria for publication, they are likely to be important assets for gLM developers and users. However, despite differences in current benchmarks’ choice of tasks and methodologies, they provide seemingly redundant insight into gLMs’ capabilities. Moving forward, it will be incumbent on the computational genomics community to develop standardized and extensible benchmarks that are widely trusted.
CONCLUDING REMARKS & FUTURE PERSPECTIVES
In an age of a vast and growing number of genomic sequences, gLMs are emerging as powerful tools to extract complex patterns useful for numerous applications, including functional constraint estimation, sequence design and transfer learning. However, they do not yet represent a magical, sudden breakthrough as the term “AI” may suggest. Instead, we view them as another useful modeling tool, much like Hidden Markov Models were when they were first introduced. Often, gLMs are claimed to be “foundation models”, a term recently invented to denote models trained on broad data that can be adapted to a wide range of downstream tasks [44]. The introduction of this new term has been criticized as the word “foundation” has the connotation of substantial improvement on downstream task performance, which is an empirical question, not an inherent property of pretrained models [130]. This criticism rings even louder in new domains such as genomics, where establishing adequate benchmarks is likely to take some time.
While earlier gLMs tend to be more or less direct adaptations from NLP models, we expect that further contextualization with deep genomics expertise will reap the highest rewards. We note that evaluating the capabilities of gLMs is challenging because metrics may be misleading, especially when over-optimized. A boon for NLP is that humans are experts in natural language and, therefore, can calibrate benchmarks to match their expertise. In genomics, however, we must rely on data and expert knowledge to falsify models. This aspect of the problem makes it especially challenging and may highlight a need for engagement with subject-matter experts and deliberate experimentation for the sake of developing benchmarks. We conclude this review with a few research directions (listed in Outstanding Questions) that we believe warrant further investigation.
@@
[style=box] Outstanding Questions
@@
    Will the scaling hypothesis hold for gLMs, and for how long? Are there really that much data available, considering that most may be non-functional?
@@ Acknowledgments
Declaration of Interests
The authors declare no competing interests.